The 6-pole,wound-rotor induction motor is excited by a 3-phase, 60 Hz source. Calculate the

frequency of the rotor current under the following conditions:

(i) at standstill

(ii) motor turning at 500 rpm in the same direction as the revolving field

(iii) motor turning at 500 rpm in the opposite direction to the revolving field

(iv) motor turning at 2000 rpm in the same direction as the revolving field


1
Expert's answer
2021-08-10T05:55:06-0400

Synchronous Speed "=\frac{120f}{P}=\frac{120\times60}{6} =1200"

slip "= \frac{N_s-N}{N}"

1) at standstill Rotor Frequency is 60 Hz

2) motor turning at 500 rpm in the same direction as the revolving field

Rotor Frequency "=sf= (\frac{N_s-N}{N})\times f =(\frac{1200-500}{500})\times 60 =84" Hz

3) motor turning at 500 rpm in the opposite direction to the revolving field

Rotor Frequency "=sf= (\frac{N_s+N}{N})\times f =(\frac{1200+500}{500})\times 60 =85" Hz

4) motor turning at 2000 rpm in the same direction as the revolving field

Rotor Frequency "=sf= (\frac{N_s-N}{N})\times f =(\frac{1200-2000}{2000})\times 60 =24" Hz